OpenStudy (anonymous):

15.15 cm^3/12.0 cm^2

OpenStudy (anonymous):

15.15^3/12.0^2 =3477.265875/144 =24.14767969 =24.15 (round to 2 decimals because you are only given 2 decimals)

OpenStudy (anonymous):

ummmmmm. my answer is 1.3cm but how

OpenStudy (anonymous):

you have to do the exponents first (BEDMAS)

OpenStudy (anonymous):

15.15cm cubed _______________ = 1.3 cm rounded up 12.0 cm squared I used a calculater then figurede cm^3 goes into cm^2 once and cm is left

OpenStudy (anonymous):

sorry wrote it wrong its centimeters cubed and squared

OpenStudy (anonymous):

12 goes into 15 once with a little left it is 1.3